OpenStudy (anonymous):

What is the critical angle of a diamond having a refractive index of 2.42? Use air as the second medium. Air has the refractive index 1.00.

OpenStudy (anonymous):

14.9 24.4 36.6 40.9

OpenStudy (anonymous):

Snell's law is \[n_1\sin\theta_1 = n_2\sin\theta_2\] Let's say material 1 is diamond and material 2 is air. The critical angle \(\theta_1 = \theta_c\) in diamond is reached when the angle in air is \(\theta_2 = 90^\circ\) Therefore \[2.42 \sin\theta_c = 1.00 \sin90^\circ = 1\] solve for \(\theta_c\)